what is the specific activity (in Ci/g) of an isotope if 3.56 mg emits 4.26 x 10^8 beta particles per second?
Answer #1

Answer

3.23 Ci/g

Explanation

Activity in Becquerel = 4.26×108 Bq

1Ci = 3.7 ×1010 Bq

(One Curie approximately the activity of one gram of radium per second)

Therefore,

4.26 × 108 Bq × ( 1Curie/3.7 ×1010Bq ) = 1.1514 × 10-2 Ci

Specific activity of the isotope = 1.1514×10-2Ci/0.00356g = 3.23 Ci/g